OpenStudy (anonymous):

Can anyone help me please? Which expression will simplify to 1? a. (m+9)/(m-9) * (m+9)/(m-9) b. (m+9)/(m-9) * (m-9)/(m+9) c. (m+9)/(m-9) * (9+m)/(9-m) d. (m+9)/(m-9) * (9-m)/(9+m)

OpenStudy (johnweldon1993):

Hint \(\Large \frac{a}{b} \times \frac{c}{d} = \frac{ac}{bd}\) If by chance we have \(\Large \frac{a}{b} \times \frac{b}{a} = \frac{ab}{ba} = \frac{\cancel{a}\cancel{b}}{\cancel{b}\cancel{a}} = 1\)

OpenStudy (anonymous):

So it would be b because '(m+9)/(m-9) * (m-9)/(m+9)' would cancel out (if i did the math correctly)

OpenStudy (anonymous):

yeah it's b

OpenStudy (anonymous):

GRACIAS AMIGOS! (Thank You 'Friends')

OpenStudy (johnweldon1993):

Not a problem!